OpenStudy (cwrw238):

T've got a mental block with this one - my cousin's homework. The first part is easy but i'm not sure how to proceed with the second. Verify that x = 2 and y = -1 satisfy the system of equations x^3 - 3xy^2 = 2, 3x^2y - y^3 = -11. Hence find the cube root of 2 - 11i.

OpenStudy (cwrw238):

Its probably obvious but just can't get my head around it somehow..

OpenStudy (math&ing001):

Well yeah it is quite easy: 2-11i = x^3 -3xy^2 + 3x^2y - y^3 = (x-y)^3

OpenStudy (amistre64):

if x=2 and y=-1 satisfy the system; then they will both be true for the stated values

OpenStudy (amistre64):

i spose this is the 'second' part: Hence find the cube root of 2 - 11i.

OpenStudy (cwrw238):

i know if i let (x + yi)^3 = 2 - 11i then plug in x = 2 and y = -1 into the above then LHS = RHS but how can i justify that?

OpenStudy (math&ing001):

It's already justified by part A.

OpenStudy (cwrw238):

so cube root = 2 - i

OpenStudy (math&ing001):

Yep !

OpenStudy (cwrw238):

yes

OpenStudy (cwrw238):

not difficult!

OpenStudy (cwrw238):

lol

OpenStudy (math&ing001):

Indeed :P